# Generalizable Prompt Learning of CLIP: A Brief Overview

### v1

@@ 4 Experimental Dataset
As shown in Table 4, the dataset covers fine-grained classification, scene recognition, action recognition, and texture classification. The ImageNet [20] dataset was constructed by Stanford University, The training set contains approximately 1.2 million images, while the validation set contains 50000 images. The Caltech101 [30] dataset covers common objects, animals, and scenes such as food, vehicles, furniture, etc. OxfordPets [74] is a pet image dataset provided by the University of Oxford. StanfordCars [45] was released by Stanford University’s Artificial Intelligence Laboratory in 2013. The Flowers102 [72] dataset was released by the University of Oxford in 2008, which includes 102 categories and a total of 8189 flower images. The Food101 [4] dataset is a commonly used fine-grained food classification dataset provided by ETH Zurich. The FGVCAircraft [69] dataset is a fine-grained aircraft classification dataset that includes 100 aircraft categories. The SUN397 [98] dataset is a large-scale scene understanding dataset provided by Princeton University. The DTD [16] dataset is a texture classification dataset provided by the University of Oxford, containing 47 texture categories. The EuroSAT [34] dataset is collected based on satellite images. The UCF101 [89] dataset is a real-world action recognition dataset provided by the University of Florida. For the domain transfer dataset, we use four variants of ImageNet [82, 93, 36, 35].
Dataset Classes Train Val Test
ImageNet [20] 1,000 1.28 M N/A 50,000
Caltech101 [30] 100 4,128 1,649 2,465
OxfordPets [74] 37 2,944 736 3,669
Flowers102 [72] 102 4,093 1,633 2,463
StanfordCars [45] 196 6,509 1,635 8,041
Food101 [4] 101 50,500 20,200 30,300
FGVCAircraft [69] 100 3,334 3,333 3,333
SUN397 [98] 397 15,880 3,970 19,850
EuroSAT [34] 10 13,500 5,400 8,100
DTD [16] 47 2,820 1,128 1,692
UCF101 [89] 101 7,639 1,898 3,783
ImageNet-V2 [82] 1,000 N/A N/A 10,000
ImageNet-Sketch [93] 1,000 N/A N/A 50,889
ImageNet-A [36] 200 N/A N/A 7,500
ImageNet-R [35] 200 N/A N/A 30,000
Table 4: Training and testing datasets. This dataset consists of $1$ image classification datasets and 4 variant datasets of ImageNet.
5 Training Implementation
@@ 7 Analysis
  • Most of the existing work is based on ViT-B/16, but there are relatively few experiments conducted on relatively advanced architectures such as EVA-CLIP [24].

### v2

@@ 4 Experimental Dataset
The dataset covers fine-grained classification, scene recognition, action recognition, and texture classification. The ImageNet [20] dataset was constructed by Stanford University, The training set contains approximately 1.2 million images, while the validation set contains 50000 images. The Caltech101 [29] dataset covers common objects, animals, and scenes such as food, vehicles, furniture, etc. OxfordPets [73] is a pet image dataset provided by the University of Oxford. StanfordCars [44] was released by Stanford University’s Artificial Intelligence Laboratory in 2013. The Flowers102 [71] dataset was released by the University of Oxford in 2008, which includes 102 categories and a total of 8189 flower images. The Food101 [4] dataset is a commonly used fine-grained food classification dataset provided by ETH Zurich. The FGVCAircraft [68] dataset is a fine-grained aircraft classification dataset that includes 100 aircraft categories. The SUN397 [97] dataset is a large-scale scene understanding dataset provided by Princeton University. The DTD [16] dataset is a texture classification dataset provided by the University of Oxford, containing 47 texture categories. The EuroSAT [33] dataset is collected based on satellite images. The UCF101 [88] dataset is a real-world action recognition dataset provided by the University of Florida. For the domain transfer dataset, we use four variants of ImageNet [81, 92, 35, 34].
5 Training